OpenStudy (anonymous):

I cant figure out how to simplify this problem please help! a^4-81 ___________ 5a-15

OpenStudy (anonymous):

that means over

OpenStudy (kirbykirby):

\[ \frac{a^4-81}{5a-15}=\frac{\color{blue}{(a^2-9)}(a^2+9)}{5(a-3)}=\frac{\color{blue}{(a-3)(a+3)}(a^2+9)}{5(a-3)}=\frac{(a+3)(a^2+9)}{5}\]

OpenStudy (anonymous):

that isnt the awnser

OpenStudy (kirbykirby):

Maybe your answer has the numerator multiplied out ?

OpenStudy (kirbykirby):

\[ \frac{a^3+3a^2+9a+27}{5}\]?